OpenStudy (anonymous):

Which property of equality would be used to isolate the variable a in this equation? a – 4 = 12 A. addition property of equality B. subtraction property of equality C. multiplication property of equality D. division property of equality

OpenStudy (anonymous):

well it would be A obviously a - 4 + 4 = 12 + 4 a = 16

OpenStudy (anonymous):

Which property of equality would be used to isolate the variable x in this equation? x – 8 = 14 would this one be that same answer too right?? @FriedRice

OpenStudy (anonymous):

yes